OpenStudy (baldymcgee6):

Scientists discover a fossil that contains 74% as much radioactive carbon, 14C, as does a living mammal. Assuming the half-life of 14 C to be about 5730 years, write the ODE representing the decay of carbon in the specimen and estimate its age.

OpenStudy (anonymous):

whats ODE

OpenStudy (baldymcgee6):

ordinary differential equation

OpenStudy (anonymous):

\[\frac{ \delta fc }{ \delta t }=-\frac{ \ln(2) }{ half-life }fc\]

OpenStudy (anonymous):

where fc is the concentration of C14

OpenStudy (baldymcgee6):

right, I know the formula for half-life but i don't know how to get the ODE

OpenStudy (anonymous):

thats the ODE

OpenStudy (anonymous):

you want the process to get there?

OpenStudy (baldymcgee6):

yes please

OpenStudy (anonymous):

Im too tired to explain how i got there, its a little bit complicated... srry mate

OpenStudy (baldymcgee6):

Alright, thanks anyways.

OpenStudy (anonymous):

but do u know how to solve that?